A person is drinking water from a plastic bottle. When the water level is 6cm from the top of the bottle, the person blows across the top producing a sound. What is the frequency of the sound produced?
Answer #1

we have wavelenght = d/4 =6/4 = 1.5 cm =0.015 m


vel of sound = freq* wavelength


freq of sound = 1482/0.015 = 98800 Hz